# Quantum metrology through spectral measurements in quantum optics

### v1

@@ Acknowledgements.
The authors are thankful to Alejandro González Tudela, Roberto Di Candia, Simone Felicetti, and Maryam Khanahmadi for insightful discussions. C.S.M. acknowledges support by the project PID2023-149969NA-I00 funded by the Spanish Agencia Estatal de Investigación MICIU/AEI/10.13039/501100011033. A.V.V. also acknowledges support from the Swedish Foundation for Strategic Research (grant number FFL21-0279).
@@ A.1 Single-sensor cascaded master equation
We now have operators evaluated at two different times, $t$ and $t-\tau$. However, since we are considering a one-way driving scenario and $\tau>0$, the effect of this retarded time is merely a time shift of the origin of the time axis for the second system. Thus, we can set $\tau\rightarrow 0^{+}$ without losing generality while maintaining causality [60]. We also note that, given the cascaded input-output relations in Eqs. 21 and 22, we can identify the input and output fields of the combined systems just as $\hat{b}_{\text{in}}\equiv\hat{b}_{\text{in},1}$ and $\hat{b}_{\text{out}}\equiv\hat{b}_{\text{out},2}$ [see Fig. 7(a)]. Finally, by considering a general operator $\hat{a}$, which may act either in the Hilbert space of the source or the target, we obtain the cascaded quantum Langevin equation:
$\frac{d\hat{a}}{dt}=-i[\hat{a},\hat{H}]-\sqrt{\Gamma_{1}\Gamma_{2}}\left([\hat{a},\hat{c}^{\dagger}_{2}]\hat{c}_{1}-\hat{c}_{1}^{\dagger}[\hat{a},c_{2}]\right)\\ +\sum_{i=1}^{2}\left[-[\hat{a},\hat{c}_{i}^{\dagger}]\left(\frac{\Gamma_{i}}{2}\hat{c}_{i}+\sqrt{\Gamma_{i}}\hat{b}_{\text{in}}\right)+\left(\frac{\Gamma_{i}}{2}\hat{c}^{\dagger}_{i}+\sqrt{\Gamma_{i}}\hat{b}^{\dagger}_{\text{in}}\right)[\hat{a},\hat{c}_{i}]\right].$ (23)
@@
All other products, including $dtd\hat{B}(t)$ or $dtd\hat{B}^{\dagger}(t)$, and higher orders are set to zero. We note that these quantum Ito rules are valid when one takes the expected value of the operators. Hence, using the quantum Ito algebra and taking the expected value of the quantum Langevin equation in (23), we obtain the cascaded master equation:
$\frac{d\rho(t)}{dt}=-i[\hat{H},\hat{\rho}]+\sum_{i}^{2}\frac{\Gamma_{i}}{2}\mathcal{D}[\hat{c}_{i}]\hat{\rho}\\ -\sqrt{\Gamma_{1}\Gamma_{2}}\left([\hat{c}_{2}^{\dagger},\hat{c}_{1}\hat{\rho}]+[\hat{\rho}\hat{c}_{1}^{\dagger},\hat{c}_{2}]\right).$ (25)
This equation contains the standard terms of a Lindblad master equation: a term for the free evolution governed by the Hamiltonian $\hat{H}$, and Lindblad terms for both the source and the target, with corresponding dissipative rates $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Additionally, it features an extra term that accounts for the asymmetric cascaded coupling between the source and the target. We can see from Eq (25) that it is exactly the same as Eq. (3) by identifying: $\hat{H}\rightarrow\hat{H}+\Delta_{\xi}\hat{\xi}^{\dagger}\hat{\xi}$, $\hat{c}_{1}\rightarrow\hat{\sigma}$, $\hat{c}_{2}\rightarrow\hat{\xi}$, $\Gamma_{1}\rightarrow\gamma$, $\Gamma_{2}\rightarrow\Gamma$, and $\sqrt{\Gamma_{1}\Gamma_{2}}\rightarrow\sqrt{\varepsilon\gamma\Gamma}$.
@@ A.2 Two-sensor cascaded master equation
$\frac{d\rho(t)}{dt}=-i[\hat{H},\hat{\rho}]+\frac{\Gamma_{1}}{2}\mathcal{D}[\hat{c}_{1}]\hat{\rho}+\sum_{\alpha=1}^{2}\frac{\Gamma_{2}^{(\alpha)}}{2}\mathcal{D}[\hat{c}_{2}^{(\alpha)}]\hat{\rho}-\sum_{\alpha=1}^{2}\sqrt{\Gamma_{1}\Gamma_{2}^{(\alpha)}}\left([\hat{c}_{2}^{(\alpha)\dagger},\hat{c}_{1}\hat{\rho}]+[\hat{\rho}\hat{c}_{1}^{\dagger},\hat{c}_{2}^{(\alpha)}]\right).$ (31)
We observe from Eq (31) that is formally the same as Eq. (4) by identifying: $\hat{H}\rightarrow\hat{H}+\Delta_{1}\hat{\xi}^{\dagger}_{1}\hat{\xi}_{1}+\Delta_{2}\hat{\xi}^{\dagger}_{2}\hat{\xi}_{2}$, $\hat{c}_{1}\rightarrow\hat{\sigma}$, $\hat{c}_{2}^{(1)}\rightarrow\hat{\xi}_{1}$, $\hat{c}_{2}^{(2)}\rightarrow\hat{\xi}_{2}$, $\Gamma_{1}\rightarrow\gamma$, $\Gamma_{2}^{(1/2)}\rightarrow\Gamma$, and $\sqrt{\Gamma_{1}\Gamma_{2}}\rightarrow\sqrt{\varepsilon\gamma\Gamma}$. This formalism captures the essential physics of cascaded coupling mediated by a beam splitter and rigorously accounts for the associated vacuum fluctuations.
APPENDIX B Method to compute the steady-state and its derivative
@@ B.2 Computation of the differentiation of the steady-state
The formal relation from (36) can be extended to compute the differentation of the steady-state density matrix with respect to a system parameter $\theta$, denoted as $\partial_{\theta}\hat{\rho}_{\text{ss}}$. This is achieved by taking the derivative with respect $\theta$ in (33), such that
@@
Note that before computing the differentiation of the steady state with respect to a parameter $\theta$, the steady state itself must first be determined. Using the formal relation for computing the steady-state in (33), we can rewrite (39) solely in terms of the effective Liouvillian,
@@
Figure 8 shows the CFI, $F_{\theta}$, as a function of the sensor-laser detuning $\Delta_{\xi}$ (top panels) and of the driving strength $\Omega$ (bottom panels), with panel (a) corresponding to $\theta=\Omega$ and panel (b) to $\theta=\gamma$. Similar to the case discussed in the main text, the CFI for these parameters exhibits rich spectral features that closely resemble the Mollow triplet structure [cf. Fig. 2(b)]. In fact, for these two particular cases, we can observe how the CFI features local minima around the Mollow resonances at $\Delta_{\xi}\approx\pm 2\Omega$. For $\theta=\Omega$ [panel (a)], we find that photon-counting measurements with $\alpha=0$ (dashed curves) nearly saturate the QFI (solid curves) across the frequency domain, as illustrated for $\Omega={\gamma/(2\sqrt{2}),3\gamma}$ in blue and red, respectively. In this case, standard photon-counting essentially constitutes the optimal measurement strategy.
@@
Figure 8: Frequency-resolved Fisher information for one-sensor to estimate (a) $\theta=\Omega$ and (b) $\theta=\gamma$. The top panels of each figure shows the classical Fisher information $F_{\theta}$ (dashed curve) and the quantum Fisher information $I_{\theta}$ (solid curves) for two values of $\Omega=\{\gamma/2\sqrt{2},3\gamma\}$, in blue and red, respectively. The lower panels depict $F_{\theta}$ in terms of the driving strength $\Omega$ and the cavity frequency $\Delta_{\xi}$. Parameters: (a) $\Omega=3\gamma$; (a-c) $\Delta=10^{-6}\gamma$, $\Gamma=10^{-1}\gamma$, $\varepsilon=0.1$.
@@
In this section, we investigate how imperfect system-sensor coupling— characterized by the parameter $\varepsilon$—affects the classical Fisher information. For clarity and simplicity, we focus on the single-sensor scenario with no displacement ($\alpha=0$) and use the expression of the CFI in terms of normally ordered correlators presented in Eq. (12).
@@
$G^{(k)}_{\Gamma}=\frac{g^{2}}{\Gamma^{k}}S^{(k)}_{\Gamma}$, (42)
@@
$G^{(k)}_{\Gamma,\varepsilon_{1}}=\left(\frac{\varepsilon_{1}}{\varepsilon_{2}}\right)^{k}G^{(k)}_{\Gamma,\varepsilon_{2}}$. (43)
Considering $\varepsilon_{2}=1$ (perfect coupling), we observe that the strength of $k$th order correlators is increasingly suppressed with decreasing $\varepsilon$, due to the exponential dependence on the order $k$. Consequently, the estimation sensitivity—quantified by the CFI—is expected to deteriorate significantly under imperfect coupling.
Figure 9: Impact of imperfect system-sensor coupling ($\varepsilon$) on the estimation sensitivity. (a) $F_{\Delta}$ as a function of the sensor-laser detuning $\Delta_{\xi}$, shown for various coupling efficiencies $\varepsilon={10^{-6},10^{-4},10^{-2},1}$, with lighter to darker curves representing increasing values of $\varepsilon$. (b) $F_{\Delta}$ as a function of both $\Delta_{\xi}$ and $\varepsilon$. Parameters: $\Delta=0$, $\Omega=10\gamma$, and $\Gamma=0.1\gamma$.
Substituting this scaling into Eq. (12), the CFI under imperfect coupling becomes:
$F^{\alpha}_{\theta,\text{imperfect}}=\sum_{n}^{n_{\text{exc}}}\frac{\left[\sum_{k\geq n}^{n_{\text{exc}}}c_{n,k}\varepsilon^{k}\partial_{\theta}G^{(k)}_{\Gamma,\text{perfect}}\right]^{2}}{\sum_{k\geq n}^{n_{\text{exc}}}c_{n,k}\varepsilon^{k}G^{(k)}_{\Gamma,\text{perfect}}}$. (44)
@@ E.2 Poissonian Fisher information bound
Considering these two conditions, Eqs. 48 and 49, we arrive at the bound
@@ F.1 General method
As explained in the main text [see inset Fig. 4(a)], the mean-field engineering technique is implemented by means of an unbalanced beam splitter combined with a local oscillator in the limit of perfect transmittance. In this scenario, the sensor receives the output field displaced by a complex amplitude $\alpha$ in phase space.
@@
where $\hat{\rho}_{\theta}$ is the density matrix of the sensor, and $\hat{D}(\alpha)=\exp[\alpha\hat{\xi}^{\dagger}-\alpha^{*}\hat{\xi}$] is the displacement operator. In fact, by projecting onto the Fock state basis via $\langle n|\cdot|n\rangle$, we obtain the generalized photon-counting POVM introduced in Eq. (7).
@@
The matrix elements of the displacement operator in the Fock basis are easiliy given by the identity [102, 103]:
@@ F.2 Isolating quantum fluctuations
The classical Fisher information for these two parameters, $\theta=\Omega$ [Fig. 10(a)] and $\theta=\gamma$ [Fig. 10(b)], reaches significantly larger values than in the case $\theta=\Delta$. Notably, standard photon-counting measurements—i.e., $\vec{\alpha}=\vec{0}$–already saturate the quantum Fisher information (QFI), indicating that no further measurement optimization is required to attain the ultimate precision limit, as we already observed for the case of a single sensor in Fig. 8 [see Appendix C.]. Moreover, when comparing the joint two-sensor CFI with its independent counterpart, $F_{\theta}^{I}$ [see Eq. (18)], we can observe that there is practically no metrological gain from retaining cross-correlations. As shown in the lower-right panels of Fig. 10, retaining cross-correlations provides at most a threefold improvement over independent measurements, in stark contrast with the case $\theta=\Delta$ (see Fig. 5), where cross-correlations yielded enhancements of up to five orders of magnitude near leapfrog resonances.
@@ H.1 Proof
while the $\mathbb{M}$ is the coefficient matrix determined by Eq. (75). However, since the coefficient matrix $\mathbb{M}$ is not upper-triangular, a direct inversion of this system is not possible, meaning that we cannot establish a clear bijection between probability distribution and correlators due to redundancies: the number of variables exceeds the number of equations. To solve this, we impose three physical constraints:
@@
On this reduced space, the coefficient matrix becomes lower-triangular, allowing for an explicit recursive solution. These $N^{2}$ variables do not contain the zeros ($n_{1},n_{2}=0$), so that now we can establish the bijection between correlator and probability distribution elements as now the the linear system is triangular:
@@
$\vec{P}^{\prime}$ $=(P_{1,1},P_{2,1},\ldots,P_{N,1},P_{1,2},P^{N,2},\ldots,P_{N,N}),$ (82b)
and $\mathbb{M}^{\prime}$ is now the coefficient matrix generated by $\vec{G}^{\prime}$ given Eq. (75). As a consequence, we can now invert the relation of this matrix by solving it recursively (backward Gaussian substitution), such that

### v2

@@ Acknowledgements.
The authors are thankful to Alejandro González Tudela, Roberto Di Candia, Simone Felicetti, Maryam Khanahmadi and Victor Rueskov Christiansen for insightful discussions. C. S. M. acknowledges support by the project PID2023-149969NA-100 funded by the Spanish Agencia Estatal de Investigación MICIU/AEI/10.13039/501100011033, and by a 2025 Leonardo Grant for Scientific Research and Cultural Creation from the BBVA Foundation. A.V.V. also acknowledges support from the Swedish Foundation for Strategic Research (grant number FFL21-0279).
@@ A.1 Single-sensor cascaded master equation
We now have operators evaluated at two different times, $t$ and $t-\tau$. However, since we are considering a one-way driving scenario and $\tau>0$, the effect of this retarded time is merely a time shift of the origin of the time axis for the second system. Thus, we can set $\tau\rightarrow 0^{+}$ without losing generality while maintaining causality [48]. We also note that, given the cascaded input-output relations in Eqs. 32 and 33, we can identify the input and output fields of the combined systems just as $\hat{b}_{\text{in}}\equiv\hat{b}_{\text{in},1}$ and $\hat{b}_{\text{out}}\equiv\hat{b}_{\text{out},2}$ [see Fig. 9(a)]. Finally, by considering a general operator $\hat{a}$, which may act either in the Hilbert space of the source or the target, we obtain the cascaded quantum Langevin equation:
$\frac{d\hat{a}}{dt}=-i[\hat{a},\hat{H}]-\sqrt{\Gamma_{1}\Gamma_{2}}\left([\hat{a},\hat{c}^{\dagger}_{2}]\hat{c}_{1}-\hat{c}_{1}^{\dagger}[\hat{a},c_{2}]\right)+\sum_{i=1}^{2}\left[-[\hat{a},\hat{c}_{i}^{\dagger}]\left(\frac{\Gamma_{i}}{2}\hat{c}_{i}+\sqrt{\Gamma_{i}}\hat{b}_{\text{in}}\right)+\left(\frac{\Gamma_{i}}{2}\hat{c}^{\dagger}_{i}+\sqrt{\Gamma_{i}}\hat{b}^{\dagger}_{\text{in}}\right)[\hat{a},\hat{c}_{i}]\right].$ (34)
@@
All other products, including $dtd\hat{B}(t)$ or $dtd\hat{B}^{\dagger}(t)$, and higher orders are set to zero. We note that these quantum Ito rules are valid when one takes the expected value of the operators. Hence, using the quantum Ito algebra and taking the expected value of the quantum Langevin equation in Eq. (34), we obtain the cascaded master equation:
$\frac{d\rho}{dt}=-i[\hat{H},\hat{\rho}]+\sum_{i}^{2}\frac{\Gamma_{i}}{2}\mathcal{D}[\hat{c}_{i}]\hat{\rho}\\ -\sqrt{\Gamma_{1}\Gamma_{2}}\left([\hat{c}_{2}^{\dagger},\hat{c}_{1}\hat{\rho}]+[\hat{\rho}\hat{c}_{1}^{\dagger},\hat{c}_{2}]\right).$ (36)
This equation contains the standard terms of a Lindblad master equation: a term for the free evolution governed by the Hamiltonian $\hat{H}$, and Lindblad terms for both the source and the target, with corresponding dissipative rates $\Gamma_{1}$ and $\Gamma_{2}$, respectively. Additionally, it features an extra term that accounts for the asymmetric cascaded coupling between the source and the target. We can see from Eq. (36) that it is exactly the same as Eq. (6) by identifying: $\hat{H}\rightarrow\hat{H}+\Delta_{\xi}\hat{\xi}^{\dagger}\hat{\xi}$, $\hat{c}_{1}\rightarrow\hat{\sigma}$, $\hat{c}_{2}\rightarrow\hat{\xi}$, $\Gamma_{1}\rightarrow\gamma$, $\Gamma_{2}\rightarrow\Gamma$, and $\sqrt{\Gamma_{1}\Gamma_{2}}\rightarrow\sqrt{\varepsilon\gamma\Gamma}$.
@@ A.2 Two-sensor cascaded master equation
$\frac{d\rho(t)}{dt}=-i[\hat{H},\hat{\rho}]+\frac{\Gamma_{1}}{2}\mathcal{D}[\hat{c}_{1}]\hat{\rho}+\sum_{\alpha=1}^{2}\frac{\Gamma_{2}^{(\alpha)}}{2}\mathcal{D}[\hat{c}_{2}^{(\alpha)}]\hat{\rho}-\sum_{\alpha=1}^{2}\sqrt{\Gamma_{1}\Gamma_{2}^{(\alpha)}/2}\left([\hat{c}_{2}^{(\alpha)\dagger},\hat{c}_{1}\hat{\rho}]+[\hat{\rho}\hat{c}_{1}^{\dagger},\hat{c}_{2}^{(\alpha)}]\right).$ (42)
We observe from Eq (42) that is formally the same as Eq. (7) by identifying: $\hat{H}\rightarrow\hat{H}+\Delta_{1}\hat{\xi}^{\dagger}_{1}\hat{\xi}_{1}+\Delta_{2}\hat{\xi}^{\dagger}_{2}\hat{\xi}_{2}$, $\hat{c}_{1}\rightarrow\hat{\sigma}$, $\hat{c}_{2}^{(1)}\rightarrow\hat{\xi}_{1}$, $\hat{c}_{2}^{(2)}\rightarrow\hat{\xi}_{2}$, $\Gamma_{1}\rightarrow\gamma$, $\Gamma_{2}^{(1/2)}\rightarrow\Gamma$, and $\sqrt{\Gamma_{1}\Gamma_{2}^{(\alpha)}/2}\rightarrow\sqrt{\varepsilon\gamma\Gamma/2}$. This formalism captures the essential physics of cascaded coupling mediated by a beam splitter and rigorously accounts for the associated vacuum fluctuations.
APPENDIX B Method to compute the steady-state and its differentiation
@@ B.2 Computation of the differentiation of the steady-state
The formal relation from Eq. (47) can be extended to compute the differentiation of the steady-state density matrix with respect to a system parameter $\theta$, denoted as $\partial_{\theta}\hat{\rho}_{\text{ss}}$. This is achieved by taking the derivative with respect $\theta$ in Eq. (44), such that
@@
Note that before computing the differentiation of the steady state with respect to a parameter $\theta$, the steady state itself must first be determined. Using the formal relation for computing the steady-state in Eq. (44), we can rewrite Eq. (50) solely in terms of the effective Liouvillian,
@@
Figure 10: Frequency-resolved Fisher information for one-sensor to estimate (a) $\theta=\Omega$ and (b) $\theta=\gamma$. The top panels of each figure shows the classical Fisher information $F_{\theta}$ (dashed curves) and the quantum Fisher information $I_{\theta}$ (solid curves) for two values of $\Omega=\{\gamma/2\sqrt{2},3\gamma\}$, in blue and red, respectively. The lower panels depict $F_{\theta}$ in terms of the driving strength $\Omega$ and the cavity frequency $\Delta_{\xi}$. Parameters: (a) $\Omega=3\gamma$; (a-c) $\Delta=10^{-6}\gamma$, $\Gamma=10^{-1}\gamma$, $\varepsilon=0.1$.
@@
Figure 10 shows the CFI, $F_{\theta}$, as a function of the sensor-laser detuning $\Delta_{\xi}$ (top panels) and of the driving strength $\Omega$ (bottom panels), with panel (a) corresponding to $\theta=\Omega$ and panel (b) to $\theta=\gamma$. Similar to the case discussed in the main text, the CFI for these parameters exhibits rich spectral features that closely resemble the Mollow triplet structure [cf. Fig. 2(b)]. In fact, for these two particular cases, we can observe how the CFI features local minima around the Mollow resonances at $\Delta_{\xi}\approx\pm 2\Omega$. For $\theta=\Omega$ [panel (a)], we find that photon-counting measurements with $\alpha=0$ (dashed curves) nearly saturate the QFI (solid curves) across the frequency domain, as illustrated for $\Omega=\{\gamma/2\sqrt{2},3\gamma\}$ in blue and red, respectively. In this case, standard photon-counting essentially constitutes the optimal measurement strategy.
@@
Figure 11: Impact of imperfect system-sensor coupling ($\varepsilon$) on the estimation sensitivity. (a) $F_{\Delta}$ as a function of the sensor-laser detuning $\Delta_{\xi}$, shown for various coupling efficiencies $\varepsilon=\{10^{-6},10^{-4},10^{-2},1\}$, with lighter to darker curves representing increasing values of $\varepsilon$. (b) $F_{\Delta}$ as a function of both $\Delta_{\xi}$ and $\varepsilon$. Parameters: $\Delta=0$, $\Omega=10\gamma$, and $\Gamma=0.1\gamma$.
In this section, we study how imperfect system-sensor coupling—characterized by the parameter $\varepsilon$—affects the classical Fisher information. For clarity and simplicity, we focus on the single-sensor scenario with no displacement ($\alpha=0$) and use the expression of the CFI in terms of normally ordered correlators presented in Eq. (16).
@@
$G^{(k)}_{\Gamma}=\frac{g^{2k}}{\Gamma^{k}}(2\pi)^{k}S^{(k)}_{\Gamma}$, (53)
@@
$G^{(k)}_{\Gamma,\varepsilon_{1}}=\left(\frac{\varepsilon_{1}}{\varepsilon_{2}}\right)^{2k}G^{(k)}_{\Gamma,\varepsilon_{2}}$. (54)
Considering $\varepsilon_{2}=1$ (perfect coupling), we observe that the strength of $k$th order correlators is increasingly suppressed with decreasing $\varepsilon_{1}$, due to the exponential dependence on the order $2k$. Consequently, the estimation sensitivity—quantified by the CFI—is expected to deteriorate significantly under imperfect coupling.
Substituting the relation Eq. (54) into Eq. (16), the CFI under imperfect coupling becomes:
$F^{\alpha}_{\theta,\text{imperfect}}=\sum_{n}^{n_{\text{exc}}}\frac{\left[\sum_{k\geq n}^{n_{\text{exc}}}c_{n,k}\varepsilon^{2k}\partial_{\theta}G^{(k)}_{\Gamma,\text{perfect}}\right]^{2}}{\sum_{k\geq n}^{n_{\text{exc}}}c_{n,k}\varepsilon^{2k}G^{(k)}_{\Gamma,\text{perfect}}}$. (55)
@@ E.2 Poissonian Fisher information bound
Considering these two conditions, Eqs. 59 and 60, we arrive at the bound
@@ F.1 General method
As explained in the main text [see Fig. 4(a)], the mean-field engineering technique is implemented by means of an unbalanced beam splitter combined with a local oscillator in the limit of perfect transmittance. In this scenario, the sensor receives the output field displaced by a complex amplitude $\alpha$ in phase space.
@@
[…]
@@
The matrix elements of the displacement operator in the Fock basis are easily given by the identity [117, 118]:
@@ F.2 Isolating quantum fluctuations
The classical Fisher information for these two parameters, $\theta=\Omega$ [Fig. 12(a)] and $\theta=\gamma$ [Fig. 12(b)], reaches significantly larger values than in the case $\theta=\Delta$. Notably, standard photon-counting measurements—i.e., $\vec{\alpha}=\vec{0}$–already saturate the quantum Fisher information (QFI), indicating that no further measurement optimization is required to attain the ultimate precision limit, as we already observed for the case of a single sensor in Fig. 10 [see Appendix C.]. Moreover, when comparing the joint two-sensor CFI with its independent counterpart, $F_{\theta}^{I}$ [see Eq. (22)], we can observe that there is practically no metrological gain from retaining cross-correlations. As shown in the lower-right panels of Fig. 12, retaining cross-correlations provides at most a threefold improvement over independent measurements, in stark contrast with the case $\theta=\Delta$ [see Fig. 5], where cross-correlations yielded enhancements of up to five orders of magnitude near leapfrog resonances.
@@ H.1 Proof
while $\mathbb{M}$ is determined by Eq. (86). However, since the coefficient matrix $\mathbb{M}$ is not upper-triangular, a direct inversion of this system is not possible, meaning that we cannot establish a clear bijection between probability distribution and correlators due to redundancies: the number of variables exceeds the number of equations. To solve this, we impose three physical constraints:
@@
On this reduced space, the coefficient matrix becomes lower-triangular, allowing for an explicit recursive solution. These $N^{2}$ variables do not contain the zeros ($n_{1},n_{2}=0$), so that now we can establish the bijection between correlator and probability distribution elements as the the linear system is triangular:
@@
$\vec{P}^{\prime}$ $=(P_{1,1},P_{2,1},\ldots,P_{N,1},P_{1,2},P_{N,2},\ldots,P_{N,N}),$ (93b)
and $\mathbb{M}^{\prime}$ is the reduced coefficient matrix generated by $\vec{G}^{\prime}$. As a consequence, we can now invert the relation of this linear system by solving it recursively (backward Gaussian substitution), such that